# AnyPPG: An ECG-Guided PPG Foundation Model Trained on Over 100,000 Hours of Recordings for Holistic Health Profiling

### v1

@@ INTRODUCTION
Out-of-clinic health monitoring plays an important role in reducing the societal burden of disease and facilitating early detection and prevention. With the rapid proliferation of wearable technologies, physiological signal-based monitoring has emerged as a scalable and cost-effective solution for continuous assessment of health beyond conventional clinical environments 1, 2, 3. Photoplethysmography (PPG), a noninvasive optical technique that measures dynamic changes in peripheral blood volume, offers a practical means to characterize cardiovascular and systemic physiological states 4, 5, 6. Recent advances in deep learning have further enhanced the utility of PPG as an accessible modality for digital health applications, including heart rate estimation 7, hypertension screening 8, atrial fibrillation detection 9, and cardiovascular risk stratification 10.
@@
Despite substantial progress in PPG analysis, most existing studies remain constrained by limited dataset sizes, scarce labeled data, and a narrow scope of downstream tasks. These limitations restrict both model performance and the broader potential of PPG for holistic health profiling, highlighting the need for a unified framework capable of learning robust and transferable representations from large-scale data, together with a systematic investigation into the full extent of PPG capabilities. Recent advances in foundation model technology provide a promising pathway to address these challenges. By pretraining on large, heterogeneous, and multi-source datasets, often through self-supervised learning, foundation models can capture intrinsic patterns within complex physiological data and adapt effectively to diverse downstream tasks 11, 12. Such models have already demonstrated transformative potential across multiple biomedical modalities, including computational pathology 13, 14, 15, echocardiography 16, polysomnography (PSG) 17, 18, and electrocardiography (ECG) 19, 20, 21, 22. Building on these advances, we aim to extend this paradigm to develop an accurate and generalizable PPG foundation model and leverage it to systematically explore the breadth of information that PPG can provide for digital health applications.
@@
In this study, we introduce AnyPPG, a foundation model developed to explore the full potential of PPG in digital health applications. AnyPPG was pretrained on over 100,000 hours of synchronized PPG and ECG recordings from 58,796 participants across five public datasets, enabling the learning of generalizable physiological representations. Unlike previous approaches that relied solely on unimodal PPG data during pretraining 23, 24, 25, AnyPPG performs cross-modal semantic alignment between PPG and ECG representations, enhancing the physiological fidelity and transferability of learned features. This design is motivated by two key considerations. First, ECG captures cardiac electrical activity that is intrinsically coupled with the hemodynamic variations reflected in PPG, providing complementary electrophysiological information that guides representation learning. Second, mounting evidence indicates that multimodal representation learning yields richer and more transferable embeddings than single-modality training 26, 27, 28.
@@
We systematically evaluated the performance of AnyPPG across a broad spectrum of health-related tasks. Using six public datasets, the model was first assessed on eleven conventional physiological analysis tasks, including heart rate estimation and atrial fibrillation detection. Recognizing that PPG reflects peripheral hemodynamics and circulatory dynamics inherently connected to multiple organ systems, we further investigated its potential for multi-organ disease diagnosis using the Multimodal Clinical Monitoring in the Emergency Department (MC-MED) dataset. Specifically, we analyzed 1,014 three-digit International Classification of Diseases, Tenth Revision (ICD-10) categories spanning chapters I-XV to comprehensively characterize the diagnostic landscape of PPG. Across all physiological analysis tasks, AnyPPG consistently outperformed the state-of-the-art model PaPaGei 23 and achieved strong diagnostic performance (area under the receiver operating characteristic curve [AUC] > 0.70) in 137 diseases. Beyond cardiovascular conditions such as heart failure and atrial fibrillation, AnyPPG demonstrated diagnostic utility across diverse non-cardiovascular domains, including Parkinson’s and Alzheimer’s diseases in the nervous system, osteoporosis and arthritis-related disorders in the musculoskeletal system, age-related cataract and glaucoma in ocular diseases, and female genital prolapse and chronic kidney disease in the genitourinary system. These findings indicate that the physiological information captured by PPG extends far beyond cardiovascular applications, underscoring its potential as a scalable biomarker for assessing health across multiple organ systems. The development of AnyPPG provides a unified framework for advancing the understanding of the physiological relevance of PPG across organ systems and lays the groundwork for future research into its broader applications in comprehensive digital health assessment.
@@ Results
The pretraining objective of AnyPPG was to align the representations of synchronized PPG and ECG signals within a shared embedding space, enabling the model to learn physiologically grounded and generalizable representations of PPG waveforms. The pretraining was performed using synchronized PPG-ECG recordings from five publicly available datasets: MC-MED 29, PulseDB 30, the Multi-Ethnic Study of Atherosclerosis (MESA) 31, the Human Sleep Project (HSP) 32, and the Cleveland Family Study (CFS) 33. In total, these datasets comprised 109,909 hours of recordings from 58,796 subjects. Detailed descriptions of the pretraining datasets are provided in Table 1. Within MC-MED, only subjects with synchronized PPG-ECG recordings were used for pretraining, whereas the remaining subjects were reserved for downstream multi-organ disease evaluation to ensure complete subject-level independence.
Model performance on conventional physiological analysis tasks was further assessed using six independent public datasets: PPG-DaLiA 34, the Cuff-Less Blood Pressure Estimation Dataset (UCI-BP) 35, the Brno University of Technology Smartphone PPG Database (BUT PPG) 36, the Gyro-Acc-PPG Dataset 37, the Wearable Stress and Affect Detection (WESAD) dataset 38, and DeepBeat 39. For multi-organ disease evaluation, AnyPPG was fine-tuned using the pretraining subset of MC-MED, and its performance was assessed exclusively on the held-out subjects. Detailed dataset descriptions are provided in Table 1.
@@ Methods
The MC-MED dataset 29 comprises 118,385 adult encounters collected at the Stanford Adult Emergency Department between 2020 and 2022. It contains continuously recorded vital signs and physiological waveforms, including PPG, ECG, and respiratory signals. In addition, the dataset provides detailed patient demographics, medical histories, clinical orders, medication records, laboratory and imaging results, and documented clinical outcomes.
@@
The PulseDB dataset 30 is a curated resource developed to benchmark cuffless blood pressure estimation methods. It contains synchronized 10-second segments of PPG, ECG, and arterial blood pressure waveforms from 5,361 subjects, along with demographic metadata such as age. PulseDB comprises two subsets: the MIMIC-III subset (n=2,423) and the VitalDB subset (n=2,938). The MIMIC-III subset includes recordings from patients admitted to the critical care units at Beth Israel Deaconess Medical Center between 2001 and 2012 40, whereas the VitalDB subset contains recordings from surgical patients undergoing routine or emergency procedures at Seoul National University Hospital, Republic of Korea 41.
@@
The MESA dataset 31 is a longitudinal cohort initiated between 2000 and 2002 to investigate the prevalence and progression of subclinical cardiovascular disease across diverse ethnic populations. Between 2010 and 2012, 2237 participants were enrolled in a Sleep Exam, which includes overnight PSG recordings comprising synchronized channels such as PPG and ECG.
@@
The HSP dataset 32 comprises 25,941 PSG recordings collected from approximately 19,492 unique patients at the Massachusetts General Hospital Sleep Laboratory as of April 1, 2023. Each recording includes a standardized set of physiological channels, with a subset containing synchronized PPG and ECG signals.
@@
The CFS dataset 33 is a longitudinal family-based cohort established to quantify the familial aggregation of sleep apnea. Initiated in 1990, the study enrolled 2,284 individuals (46% African American) from 361 families, with participants assessed up to four times over a 16-year period.
@@
All PPG and ECG recordings were preprocessed using a unified pipeline designed to generate temporally aligned and noise-suppressed inputs for pre-training. First, each continuous recording was divided into non-overlapping segments of 10 s duration, and segments containing more than 25% invalid or motionless signals were excluded. The retained segments were then band-pass filtered to remove baseline drift and high-frequency noise: PPG signals were filtered within the range of 0.5-8 Hz following the method of Elgendi et al. 42, while ECG signals were filtered within 0.5-40 Hz and further denoised with a 50 Hz notch filter to eliminate powerline interference. ECG polarity inversion was automatically detected and corrected to ensure consistent waveform morphology. Signal quality was subsequently evaluated for the ECG signals, with quality indices computed according to the criterion proposed by Zhao et al. 43. To enhance the robustness of AnyPPG to noise, no explicit quality screening was applied to the PPG signals. Afterward, the remaining segments were resampled to a uniform sampling rate of 125 Hz and standardized using z-score normalization along the temporal dimension. The resulting synchronized and normalized PPG-ECG segment pairs served as inputs for contrastive pre-training.
@@
The AnyPPG model is built upon the Net1D architecture 44, a one-dimensional convolutional neural network derived from the ResNet framework 45, 46. It consists of two structurally identical encoders that process synchronized PPG and ECG signals to extract modality-specific feature representations. The complete architecture of the encoder and the detailed configuration of its internal modules are summarized in Table 5 and Table 6. Each encoder contains approximately 5.8 million parameters, and transforms an input signal into a 1024-dimensional embedding. Subsequently, two identical projectors map the PPG and ECG embeddings into a shared embedding space. Each projector is composed of a Linear layer, a GELU activation, and another Linear layer, which sequentially reduce the feature dimension from 1024 to 512 and finally to 256. In this 256-dimensional shared space, the embeddings from both modalities are aligned through a CLIP-style contrastive learning framework 26.
@@
Pre-training was conducted on four NVIDIA H20 GPUs, with each GPU processing a batch size of 1536. The model was optimized using the AdamW optimizer 47, with an initial learning rate of $5\times 10^{-4}$, a weight decay of $1\times 10^{-2}$, and a cosine learning rate schedule. Training was performed for a total of 200,000 steps, including 40,000 warm-up steps. Gradient clipping with a maximum norm of 1.0 was applied to stabilize optimization. Both the PPG and ECG encoders were randomly initialized and jointly optimized from scratch. The final checkpoint corresponding to the best validation contrastive loss was selected.
@@
Downstream evaluation was performed across six publicly available datasets, including PPG-DaLiA 34, UCI-BP 35, BUT PPG 36, the Gyro-Acc-PPG dataset 37, WESAD 38, and DeepBeat 39. These datasets collectively cover eleven dataset-task pairs, encompassing heart rate estimation, systolic and diastolic blood pressure estimation, signal quality assessment, stress recognition (binary), affect recognition (multi-class, four levels), and atrial fibrillation detection. In addition, downstream evaluation for multi-organ disease diagnosis was conducted using the MC-MED dataset. For this task, AnyPPG was fine-tuned using only the subset of MC-MED data that overlapped with the pre-training corpus, while the evaluation strictly relied on subjects and recordings that were entirely unseen during pre-training (including all train, validation, and test partitions) to ensure subject-level independence, with 20 PPG segments randomly sampled from each hospitalization record.
@@ Discussion
In this study, we propose AnyPPG, a PPG foundation model pretrained on large-scale, multi-source synchronized PPG-ECG data. By aligning PPG and ECG representations within a shared physiological embedding space, AnyPPG learns physiologically meaningful and transferable representations of PPG signals. Unlike previous foundation models such as PaPaGei 23, GPT-PPG 24, and PulsePPG 25, which relied solely on unimodal PPG data during pretraining, AnyPPG leverages cross-modal physiological alignment to enhance its capacity for modeling cardiovascular dynamics and improving representation generalizability. Building on this foundation, we systematically evaluated AnyPPG across a range of downstream tasks, encompassing both conventional physiological analyses (e.g., heart rate and blood pressure estimation) and broader explorations of multi-organ disease diagnosis, to comprehensively uncover the potential of PPG in health monitoring and assessment.
@@
In the multi-organ disease diagnosis tasks, AnyPPG demonstrated strong cross-system diagnostic capability. Across 1,014 three-digit ICD-10 codes from Chapters I-XV, AnyPPG achieved an AUC greater than 0.8 for 13 diseases and greater than 0.7 for 133 diseases. Overall, PPG exhibited the highest diagnostic performance for circulatory system disorders (Figure 2(b)), accurately identifying conditions such as heart failure, valvular diseases, arrhythmias, conduction abnormalities, and hypertension (Figure 2(a)), underscoring the close physiological coupling between PPG signals and cardiovascular health. Beyond the circulatory system, AnyPPG also demonstrated substantial diagnostic potential across multiple other ICD chapters. Representative examples include neurological disorders (e.g., Parkinson’s disease, Alzheimer’s disease), genitourinary diseases (e.g., chronic kidney disease), endocrine and metabolic disorders (e.g., type 2 diabetes), musculoskeletal and connective tissue disorders (e.g., osteoporosis, arthritis), respiratory diseases (e.g., emphysema, pulmonary edema), neoplasms (e.g., monocytic leukemia, carcinoma in situ of the skin), and eye diseases (e.g., senile cataract, glaucoma). All of these conditions achieved an AUC of at least 0.73, indicating that the hemodynamic information captured by PPG signals reflects physiological states across multiple organ systems and supports cross-organ disease recognition. Importantly, several of these diagnostic findings are physiologically interpretable. For example, chronic kidney disease and diabetes frequently co-occur within the cardiovascular-kidney-metabolic (CKM) syndrome framework 48, 49, which is characterized by intertwined cardiovascular, renal, and metabolic dysfunctions. Given that PPG signals reflect peripheral hemodynamic dynamics, they are well suited to capture circulatory abnormalities associated with these CKM-related conditions. Collectively, these findings support the capacity of AnyPPG to reflect systemic health status and underscore the promise of PPG-based modeling for precision health monitoring and early disease detection.

### v2

@@ INTRODUCTION
Out-of-clinic health monitoring plays an important role in reducing the societal burden of disease and facilitating early detection and prevention. With the rapid proliferation of wearable technologies, physiological signal-based monitoring has emerged as a scalable and cost-effective solution for continuous assessment of health beyond conventional clinical environments spatz2024wearable, williams2023wearable, pedroso2025leveraging. Photoplethysmography (PPG), a noninvasive optical technique that measures dynamic changes in peripheral blood volume, offers a practical means to characterize cardiovascular and systemic physiological states bayoumy2021smart, nie2024review, park2022photoplethysmogram. Recent advances in deep learning have further enhanced the utility of PPG as an accessible modality for digital health applications, including heart rate estimation pankaj2022review, hypertension screening elgendi2019use, atrial fibrillation detection pereira2020photoplethysmography, and cardiovascular risk stratification nie2025artificial.
@@
Despite substantial progress in PPG analysis, most existing studies remain constrained by limited dataset sizes, scarce labeled data, and a narrow scope of downstream tasks. These limitations restrict both model performance and the broader potential of PPG for holistic health profiling, highlighting the need for a unified framework capable of learning robust and transferable representations from large-scale data, together with a systematic investigation into the full extent of PPG capabilities. Recent advances in foundation model technology provide a promising pathway to address these challenges. By pretraining on large, heterogeneous, and multi-source datasets, often through self-supervised learning, foundation models can capture intrinsic patterns within complex physiological data and adapt effectively to diverse downstream tasks he2024foundation, moor2023foundation. Such models have already demonstrated transformative potential across multiple biomedical modalities, including computational pathology xiang2025vision, vorontsov2024foundation, lu2024multimodal, echocardiography christensen2024vision, polysomnography (PSG) thapa2024sleepfm, nie2025low, and electrocardiography (ECG) li2025electrocardiogram, zhang2025ecgfm, huang2025combining, xiao2025anyecg. Building on these advances, we aim to extend this paradigm to develop an accurate and generalizable PPG foundation model and leverage it to systematically explore the breadth of information that PPG can provide for digital health applications.
@@
In this study, we introduce AnyPPG, a foundation model developed to explore the full potential of PPG in digital health applications. AnyPPG was pretrained on over 100,000 hours of synchronized PPG and ECG recordings from 58,796 participants across five public datasets, enabling the learning of generalizable physiological representations. Unlike previous approaches that relied solely on unimodal PPG data during pretraining pillaipapagei, chen2025gpt, saha2025pulse, AnyPPG performs cross-modal semantic alignment between PPG and ECG representations, enhancing the physiological fidelity and transferability of learned features. This design is motivated by two key considerations. First, ECG captures cardiac electrical activity that is intrinsically coupled with the hemodynamic variations reflected in PPG, providing complementary electrophysiological information that guides representation learning. Second, mounting evidence indicates that multimodal representation learning yields richer and more transferable embeddings than single-modality training radford2021learning, yu2022coca, wangtoken.
@@
We systematically evaluated the performance of AnyPPG across a broad spectrum of health-related tasks. Using six public datasets, the model was first assessed on eleven conventional physiological analysis tasks, including heart rate estimation and atrial fibrillation detection. Recognizing that PPG reflects peripheral hemodynamics and circulatory dynamics inherently connected to multiple organ systems, we further investigated its potential for multi-organ disease diagnosis using the Multimodal Clinical Monitoring in the Emergency Department (MC-MED) dataset. Specifically, we analyzed 1,014 three-digit International Classification of Diseases, Tenth Revision (ICD-10) categories spanning chapters I-XV to comprehensively characterize the diagnostic landscape of PPG. Across all physiological analysis tasks, AnyPPG consistently outperformed the state-of-the-art model PaPaGei pillaipapagei and achieved strong diagnostic performance (area under the receiver operating characteristic curve [AUC] > 0.70) in 137 diseases. Beyond cardiovascular conditions such as heart failure and atrial fibrillation, AnyPPG demonstrated diagnostic utility across diverse non-cardiovascular domains, including Parkinson’s and Alzheimer’s diseases in the nervous system, osteoporosis and arthritis-related disorders in the musculoskeletal system, age-related cataract and glaucoma in ocular diseases, and female genital prolapse and chronic kidney disease in the genitourinary system. These findings indicate that the physiological information captured by PPG extends far beyond cardiovascular applications, underscoring its potential as a scalable biomarker for assessing health across multiple organ systems. The development of AnyPPG provides a unified framework for advancing the understanding of the physiological relevance of PPG across organ systems and lays the groundwork for future research into its broader applications in comprehensive digital health assessment.
@@ Results
The pretraining objective of AnyPPG was to align the representations of synchronized PPG and ECG signals within a shared embedding space, enabling the model to learn physiologically grounded and generalizable representations of PPG waveforms. The pretraining was performed using synchronized PPG-ECG recordings from five publicly available datasets: MC-MED kansal2025mc, PulseDB wang2023pulsedb, the Multi-Ethnic Study of Atherosclerosis (MESA) chen2015racial, the Human Sleep Project (HSP) sun2023human, and the Cleveland Family Study (CFS) redline1995familial. In total, these datasets comprised 109,909 hours of recordings from 58,796 subjects. Detailed descriptions of the pretraining datasets are provided in Table 1. Within MC-MED, only subjects with synchronized PPG-ECG recordings were used for pretraining, whereas the remaining subjects were reserved for downstream multi-organ disease evaluation to ensure complete subject-level independence.
Model performance on conventional physiological analysis tasks was further assessed using six independent public datasets: PPG-DaLiA reiss2019ppgdalia, the Cuff-Less Blood Pressure Estimation Dataset (UCI-BP) kachuee2015cuffless, the Brno University of Technology Smartphone PPG Database (BUT PPG) nemcova2021brno, the Gyro-Acc-PPG Dataset lee2018motion, the Wearable Stress and Affect Detection (WESAD) dataset schmidt2018introducing, and DeepBeat torres2020multi. For multi-organ disease evaluation, AnyPPG was fine-tuned using the pretraining subset of MC-MED, and its performance was assessed exclusively on the held-out subjects. Detailed dataset descriptions are provided in Table 1.
@@ Methods
The MC-MED dataset kansal2025mc comprises 118,385 adult encounters collected at the Stanford Adult Emergency Department between 2020 and 2022. It contains continuously recorded vital signs and physiological waveforms, including PPG, ECG, and respiratory signals. In addition, the dataset provides detailed patient demographics, medical histories, clinical orders, medication records, laboratory and imaging results, and documented clinical outcomes.
@@
The PulseDB dataset wang2023pulsedb is a curated resource developed to benchmark cuffless blood pressure estimation methods. It contains synchronized 10-second segments of PPG, ECG, and arterial blood pressure waveforms from 5,361 subjects, along with demographic metadata such as age. PulseDB comprises two subsets: the MIMIC-III subset (n=2,423) and the VitalDB subset (n=2,938). The MIMIC-III subset includes recordings from patients admitted to the critical care units at Beth Israel Deaconess Medical Center between 2001 and 2012 johnson2016mimic, whereas the VitalDB subset contains recordings from surgical patients undergoing routine or emergency procedures at Seoul National University Hospital, Republic of Korea lee2022vitaldb.
@@
The MESA dataset chen2015racial is a longitudinal cohort initiated between 2000 and 2002 to investigate the prevalence and progression of subclinical cardiovascular disease across diverse ethnic populations. Between 2010 and 2012, 2237 participants were enrolled in a Sleep Exam, which includes overnight PSG recordings comprising synchronized channels such as PPG and ECG.
@@
The HSP dataset sun2023human comprises 25,941 PSG recordings collected from approximately 19,492 unique patients at the Massachusetts General Hospital Sleep Laboratory as of April 1, 2023. Each recording includes a standardized set of physiological channels, with a subset containing synchronized PPG and ECG signals.
@@
The CFS dataset redline1995familial is a longitudinal family-based cohort established to quantify the familial aggregation of sleep apnea. Initiated in 1990, the study enrolled 2,284 individuals (46% African American) from 361 families, with participants assessed up to four times over a 16-year period.
@@
All PPG and ECG recordings were preprocessed using a unified pipeline designed to generate temporally aligned and noise-suppressed inputs for pre-training. First, each continuous recording was divided into non-overlapping segments of 10 s duration, and segments containing more than 25% invalid or motionless signals were excluded. The retained segments were then band-pass filtered to remove baseline drift and high-frequency noise: PPG signals were filtered within the range of 0.5-8 Hz following the method of Elgendi et al. elgendi2012analysis, while ECG signals were filtered within 0.5-40 Hz and further denoised with a 50 Hz notch filter to eliminate powerline interference. ECG polarity inversion was automatically detected and corrected to ensure consistent waveform morphology. Signal quality was subsequently evaluated for the ECG signals, with quality indices computed according to the criterion proposed by Zhao et al. zhao2018sqi. To enhance the robustness of AnyPPG to noise, no explicit quality screening was applied to the PPG signals. Afterward, the remaining segments were resampled to a uniform sampling rate of 125 Hz and standardized using z-score normalization along the temporal dimension. The resulting synchronized and normalized PPG-ECG segment pairs served as inputs for contrastive pre-training.
@@
The AnyPPG model is built upon the Net1D architecture hong2020holmes, a one-dimensional convolutional neural network derived from the ResNet framework he2016deep, hu2018squeeze. It consists of two structurally identical encoders that process synchronized PPG and ECG signals to extract modality-specific feature representations. The complete architecture of the encoder and the detailed configuration of its internal modules are summarized in Table 5 and Table 6. Each encoder contains approximately 5.8 million parameters, and transforms an input signal into a 1024-dimensional embedding. Subsequently, two identical projectors map the PPG and ECG embeddings into a shared embedding space. Each projector is composed of a Linear layer, a GELU activation, and another Linear layer, which sequentially reduce the feature dimension from 1024 to 512 and finally to 256. In this 256-dimensional shared space, the embeddings from both modalities are aligned through a CLIP-style contrastive learning framework radford2021learning.
@@
Pre-training was conducted on four NVIDIA H20 GPUs, with each GPU processing a batch size of 1536. The model was optimized using the AdamW optimizer loshchilov2017fixing, with an initial learning rate of $5\times 10^{-4}$, a weight decay of $1\times 10^{-2}$, and a cosine learning rate schedule. Training was performed for a total of 200,000 steps, including 40,000 warm-up steps. Gradient clipping with a maximum norm of 1.0 was applied to stabilize optimization. Both the PPG and ECG encoders were randomly initialized and jointly optimized from scratch. The final checkpoint corresponding to the best validation contrastive loss was selected.
@@
Downstream evaluation was performed across six publicly available datasets, including PPG-DaLiA reiss2019ppgdalia, UCI-BP kachuee2015cuffless, BUT PPG nemcova2021brno, the Gyro-Acc-PPG dataset lee2018motion, WESAD schmidt2018introducing, and DeepBeat torres2020multi. These datasets collectively cover eleven dataset-task pairs, encompassing heart rate estimation, systolic and diastolic blood pressure estimation, signal quality assessment, stress recognition (binary), affect recognition (multi-class, four levels), and atrial fibrillation detection. In addition, downstream evaluation for multi-organ disease diagnosis was conducted using the MC-MED dataset. For this task, AnyPPG was fine-tuned using only the subset of MC-MED data that overlapped with the pre-training corpus, while the evaluation strictly relied on subjects and recordings that were entirely unseen during pre-training (including all train, validation, and test partitions) to ensure subject-level independence, with 20 PPG segments randomly sampled from each hospitalization record.
@@ Discussion
In this study, we propose AnyPPG, a PPG foundation model pretrained on large-scale, multi-source synchronized PPG-ECG data. By aligning PPG and ECG representations within a shared physiological embedding space, AnyPPG learns physiologically meaningful and transferable representations of PPG signals. Unlike previous foundation models such as PaPaGei pillaipapagei, GPT-PPG chen2025gpt, and PulsePPG saha2025pulse, which relied solely on unimodal PPG data during pretraining, AnyPPG leverages cross-modal physiological alignment to enhance its capacity for modeling cardiovascular dynamics and improving representation generalizability. Building on this foundation, we systematically evaluated AnyPPG across a range of downstream tasks, encompassing both conventional physiological analyses (e.g., heart rate and blood pressure estimation) and broader explorations of multi-organ disease diagnosis, to comprehensively uncover the potential of PPG in health monitoring and assessment.
@@
In the multi-organ disease diagnosis tasks, AnyPPG demonstrated strong cross-system diagnostic capability. Across 1,014 three-digit ICD-10 codes from Chapters I-XV, AnyPPG achieved an AUC greater than 0.8 for 13 diseases and greater than 0.7 for 133 diseases. Overall, PPG exhibited the highest diagnostic performance for circulatory system disorders (Figure 2(b)), accurately identifying conditions such as heart failure, valvular diseases, arrhythmias, conduction abnormalities, and hypertension (Figure 2(a)), underscoring the close physiological coupling between PPG signals and cardiovascular health. Beyond the circulatory system, AnyPPG also demonstrated substantial diagnostic potential across multiple other ICD chapters. Representative examples include neurological disorders (e.g., Parkinson’s disease, Alzheimer’s disease), genitourinary diseases (e.g., chronic kidney disease), endocrine and metabolic disorders (e.g., type 2 diabetes), musculoskeletal and connective tissue disorders (e.g., osteoporosis, arthritis), respiratory diseases (e.g., emphysema, pulmonary edema), neoplasms (e.g., monocytic leukemia, carcinoma in situ of the skin), and eye diseases (e.g., senile cataract, glaucoma). All of these conditions achieved an AUC of at least 0.73, indicating that the hemodynamic information captured by PPG signals reflects physiological states across multiple organ systems and supports cross-organ disease recognition. Importantly, several of these diagnostic findings are physiologically interpretable. For example, chronic kidney disease and diabetes frequently co-occur within the cardiovascular-kidney-metabolic (CKM) syndrome framework ndumele2023synopsis, ndumele2023cardiovascular, which is characterized by intertwined cardiovascular, renal, and metabolic dysfunctions. Given that PPG signals reflect peripheral hemodynamic dynamics, they are well suited to capture circulatory abnormalities associated with these CKM-related conditions. Collectively, these findings support the capacity of AnyPPG to reflect systemic health status and underscore the promise of PPG-based modeling for precision health monitoring and early disease detection.